A sample of gas has a volume of 2.8 L at a temperature of 27 °C. What temperature is

needed to expand the volume to 15 L? The pressure and number of moles are constant.


1
Expert's answer
2021-10-14T04:03:43-0400

"V_1 = 2.8 \;L \\

T_1 = 27 \;°C = 300 \;K\\

V_2 = 15 \;L \\

P= constant \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

T_2 = \frac{T_1 \times V_2}{V_1} \\

= \frac{300 \times 15}{2.8} \\

= 1607.1 \;K \\

= 1334 \;°C"

Answer: 1334 °C